A particular fruit's weights are normally distributed, with a mean of 519 grams and a standard deviation of 29 grams. If you pick 7 fruits at random, then 5% of the time, their mean weight will be greater than how many grams? Give your answer to the nearest gram.
Answer #1

Pgno ① Answer:- ننفخنا Given that, Mean (4) = 519 grams Standard deviation (-)-29 grams Sample size (n)= 7 5 To a = a in = 10Pgno @ a=zx at la 51.6449.X 10.96096 +519 = 18.0296 +519 - 537.02 = 537 i să = 537 grams O-